### v1

# Leveraging Auxiliary Task Relevance for Enhanced Industrial Fault Diagnosis through Curriculum Meta-learning

Abstract
The accurate diagnosis of machine breakdowns is crucial for maintaining operational safety in smart manufacturing. Despite the promise shown by deep learning in automating fault identification, the scarcity of labeled training data, particularly for equipment failure instances, poses a significant challenge. This limitation hampers the development of robust classification models. Existing methods like model-agnostic meta-learning (MAML) do not adequately address variable working conditions, affecting knowledge transfer. To address these challenges, a Related Task Aware Curriculum Meta-learning (RT-ACM) enhanced fault diagnosis framework is proposed in this paper, inspired by human cognitive learning processes. RT-ACM improves training by considering the relevance of auxiliary working conditions, adhering to the principle of “paying more attention to more relevant knowledge”, and focusing on “easier first, harder later” curriculum sampling. This approach aids the meta-learner in achieving a superior convergence state. Extensive experiments on two real-world datasets demonstrate the superiority of RT-ACM framework.
Index Terms:
Fault diagnosis, meta-learning, few-shot, curriculum, task relevance.
I Introduction
Fault diagnosis is crucial for ensuring the safe and efficient operation of machines [1]. Recently, methods based on deep learning (DL) have demonstrated considerable promise in accurately diagnosing machine faults [2]. An advanced ensemble convolutional neural network (CNN) has been introduced for this purpose, utilizing diverse local minima to enhance model adaptability [3]. Furthermore, a capsule network that builds upon CNNs has been proposed for fault classification in bearings [4]. This network combines an inception block and a regression branch, thereby increasing the model’s versatility. However, a notable challenge faced by DL methods is the requirement for extensive training data. Accumulating such data in real-world scenarios is often not feasible due to associated costs and logistical challenges [5]. When machines exhibit faults, they are typically partially (e.g., showed down) or fully (e.g., turned off) deactivated for safety reasons, resulting in a scarcity and variety of fault data [6]. Although intentional fault inductions on lab-based machines might provide data, such interventions in operational settings are strictly prohibited. This situation intensifies the challenge of developing reliable DL classifiers for fault diagnosis in real-world contexts [7].
Three key strategies have emerged to tackle the challenge of data scarcity in creating reliable fault diagnosis models: Data Augmentation (DA) [8], Transfer Learning (TL) [9], and Few-Shot Learning (FSL) approaches [10]. At its essence, DA aims to expand and diversify datasets by producing additional data from the limited original samples available. For example, data enhancement techniques that incorporate Gaussian noise, masking noise, and signal translation have been explored in [8]. Moreover, generative models, including variational autoencoders and generative adversarial networks, have been employed to generate fault data for training DL-based diagnostic models in [11]. However, a notable limitation of DA is the potential degradation in data quality and precision, especially when original datasets are sparse. This might reduce the effectiveness of models trained on such data.
To circumvent the limitations of DA, the use of prior knowledge from related tasks has been explored, leading to the concept of Transfer Learning (TL). TL methods strive to utilize knowledge derived from one task to benefit a different, albeit related, task. For instance, a comprehensive one-dimensional CNN architecture, supplemented with fine-tuning mechanisms specifically for fault diagnosis, has been presented[1]. Similarly, a CNN-based model for diagnosing faults in bearings has been proposed, with its basis being knowledge garnered from laboratory bearings data[12]. However, a potential drawback of TL techniques is their inherent emphasis on adapting models to target tasks, which may hinder their broad applicability, especially in environments with limited data [13]. Due to this constraint, there exists a growing need to enhance the overall generalization capabilities of TL methods for more effective fault diagnosis in situations characterized by data sparsity.
Few-shot learning (FSL) stands out for its capability to generalize across tasks, not merely adapting to one. It achieves this by leveraging insights from various auxiliary tasks, a method commonly known as meta-learning [14]. This stands in contrast to transfer learning, where the primary focus is adaptation to a distinct task. FSL delves deeper by leveraging insights from associated tasks to benefit the primary FSL task [15]. In practical settings, machines typically operate under varied conditions. Gathering substantial data for each specific condition is both expensive and, in some cases, impractical. Yet, when data from varied operational conditions are utilized, a multitude of auxiliary tasks can be formulated. This inherent flexibility renders meta-learning especially fitting and potent for few-shot fault diagnosis scenarios. Such FSL models specifically designed for fault diagnosis have been proposed in [16]. These models frequently incorporate metric-based meta-learning methods, with Siamese neural networks and matching networks being prominent examples.
Despite the potential of metric-based meta-learning methods for feature extraction, their efficacy is hindered by data sparsity and variations in auxiliary task distributions. Model-Agnostic Meta-Learning (MAML) [17], a optimization-based meta-learning method, diverges from traditional metric-based approaches by emphasizing the development of models with robust generalization capabilities. MAML aims to find an initial parameters set, then the model could achieve optimal performance on a series of new tasks after one or more gradient updates, which are calculated using a small amount of data from the series of tasks. Through minimal fine-tuning with target task data, MAML demonstrates superior classification performance in few-shot bearing fault diagnosis. Zhang et al. [18] proposed a MAML-based few-shot learning framework for bearing fault diagnosis with limited data. They achieved this by developing a fault diagnosis model using an improved meta-relation network. Zhou et al. [19] introduced an augmented method that integrates prior knowledge into MAML to fully utilize the prior knowledge present in both training and testing sets. Chen et al. [10] proposed a MAML based model featuring a four-layer CNN capable of feature extraction and rapid adaptation, enabling it to quickly adapt to fault diagnosis tasks under various working conditions. Although these studies have achieved some performance improvements and emphasized that MAML learns to learn, they have overlooked the difficulty level of each task and have not considered the human learning experience of easy-to-hard curriculum learning. Chen et al. [20] propose a novel curriculum hardness aware meta-Learning framework, which incorporates hard sample mining and curriculum learning into a meta-learning paradigm. Hu et al. [6] utilized a specialized task-sequencing strategy for few-shot fault diagnosis, introducing a method where auxiliary tasks are arranged from simple to complex to improve knowledge adaptability during meta-learning. Nevertheless, these existing studies have not considered the relevance between different tasks and the target task. Irrelevant auxiliary tasks may negatively impact the target task.
To simultaneously consider the human easy-to-hard curriculum learning processing and the relevance between different tasks and the target task, we propose a novel Related Task Aware Curriculum Meta-learning (RT-ACM) enhanced fault diagnosis framework, which delicately considers the relevance of patterns among different tasks within the meta-learning paradigm to alleviate data sparsity and task diversity issues. To begin with, RT-ACM extends MAML to facilitate knowledge transfer to the target task under conditions of limited data. With fault diagnosis in different working conditions treated as single tasks, the initial model weights are meta-learned for faster adaptation to limited training data in target tasks. Furthermore, we design two components in the meta-training process: related task-aware meta-learning and a curriculum strategy to explicitly address the task diversity issue.
Firstly, we exploit the heuristic idea of “pay more attention to more relevant knowledge” to enhance the meta-learner by considering the relevance of auxiliary tasks as related task-aware meta-learning [21]. This approach compels the model to learn more from a combination of tasks that are more similar and are expected to be highly diverse, thereby contributing more to generalization. Specifically, we divide each meta-training round into two stages. In each stage, we first update the meta-learner in a MAML-like manner and then conditionally re-sample a new batch of tasks, focusing on the most complex working conditions based on validation score. Secondly, we propose drawing upon a curriculum strategy to improve the convergence rate and generalization capacity of the meta-learner given the high diversity among different tasks. The underlying concept is to present training tasks to the meta-learner in an “easier first, harder later” paradigm as determined by a teacher recommender [20, 22]. This aims to guide the meta-learner towards a better state. Specifically, we pre-train a teacher recommender for each auxiliary task, use the best validation score to gauge task difficulty, and utilize this knowledge to design a curriculum for the task sampling pool at each meta-training step.
Overall, our main contributions lie in three folds:
  1. 1.
    To the best of our knowledge, it is one of the early studies that examines to what extent knowledge can be transferred from auxiliary tasks to the target task by differentiating the relevance of working conditions.
  2. 2.
    A novel Related Task Aware Curriculum Meta-learning (RT-ACM) enhanced fault diagnosis framework is proposed, which leverages both transferred knowledge and the diversity of tasks within the target task to mitigate the issue of data sparsity.
  3. 3.
    Extensive experiments have been conducted on two real-world datasets to validate the superiority of RT-ACM against state-of-the-art approaches.
The rest of this article is organized as follows. The proposed method is detailed in Section II. In Section III, the effectiveness of RT-ACM is evaluated with various case studies. Finally, Section IV concludes this article and presents future work.
II Proposed RT-ACM
Assume that $A$ auxiliary tasks are available and denoted by $\left\{T^{\mu}\right\}_{\mu=1}^{A}$. In each auxiliary task, $T^{\mu}=\left\{(x^{\mu}_{i},y^{\mu}_{i})\right\}^{M}_{i=1}$, where $x^{\mu}_{i}\in\mathcal{R}^{1\times D}$ is $i$th sample of the input data. In the field of bearing fault diagnosis, 1-D vibration signal is commonly used as the input data. The vibration signal is segmented using a sliding window to generate samples, and $D$ is the length of the sliding window. $y_{i}^{\mu}$ is the health status of $x_{i}^{\mu}$. $M$ denotes the number of samples contained in $T^{\mu}$.
A target task is denote by $T^{tar}=\left\{(x^{tar}_{i},y^{tar}_{i})\right\}^{M^{tar}}_{i=1}$, where only a few samples are available for training. Thus, $M^{tar}$, the number of samples contained in $T^{tar}$, is assigned a very small value. $x^{tar}_{i}\in\mathcal{R}^{1\times D}$ is the input data of the $i$th sample and $y^{tar}_{i}$ is the corresponding label.
The overview of RT-ACM is outlined in Fig. 1, mainly composed of two channels (i.e., meta-level training on auxiliary tasks and target-level fine-tune). In particular, the meta-level training exploits meta-learning to capture the common vibration pattern by holding the principle of “paying more attention to more relevant knowledge” and learning knowledge curriculum judge by teacher recommender. The goal of target-level fine-tune is to learn the accurate vibration pattern in the target task and performs the predictions.
Figure 1: The overall framework of our proposed RT-ACM.
II-A Meta-level Training on Auxiliary Tasks
To distil knowledge from auxiliary tasks and employ vibration patterns, we extend the MAML with long short-term memory (LSTM) [23] as the framework for the meta-learning update. In particular, we devise a related task aware meta-learning and curriculum strategy that can transfer knowledge based on the relevance of patterns among tasks.
II-A1 Basic Theory of MAML
Unlike traditional deep learning, as a category of meta-learning, the MAML algorithm is designed to train the initialization parameters of a model within the outer loop. By learning a set of initialization parameters that are sufficiently general, the model can achieve robust performance efficiently with only a few gradient update steps when adapting to new tasks with limited data. From a learning perspective, MAML seeks to identify general internal representations that facilitate transferability to various tasks. Consequently, when confronted with new tasks, only minimal parameter fine-tuning is necessary.
Following [20], the fault diagnosis in different working conditions can be viewed as a single task in a meta-learning paradigm. Thus, the vibration signal of auxiliary tasks $\left\{T^{\mu}\right\}_{\mu=1}^{A}$ are denoted as $\mathbb{D}_{meta}^{(aux)}$ and the vibration signal of target tasks $T^{tar}$ are divided as training sets $\mathbb{D}_{train}^{(tar)}$ and test sets $\mathbb{D}_{test}^{(tar)}$. We treat each working conditions $y_{m}$ as a meta-learning task, where each task has support set $\mathcal{D}^{spt}_{y_{m}}$ for training and a query set $\mathcal{D}^{qry}_{y_{m}}$ for testing. Finally, our goal is to leverage the auxiliary tasks to learn a meta-learner $F_{w}$, which contains knowledge about different working conditions.
Specifically, each iteration of MAML includes local update and global update on the sampled task batch, where the first phase updates $\theta$ locally on the $\mathcal{D}^{spt}$ of each task, and the second phase globally updates $\theta$ by gradient descent to minimize the sum of loss on the $\mathcal{D}^{qry}$ of all tasks.
  • Local update: we first sample a batch of tasks from $\mathbb{D}^{(aux)}_{meta}$. Then we randomly sample a group of vibration signal from $\mathcal{D}^{spt}_{y_{m}}$ and $\mathcal{D}^{qry}_{y_{m}}$ for each sampled tasks. Thus, we calculate the training loss on $\mathcal{D}^{spt}_{y_{m}}$ and locally update $\theta$ by one step:
    $\theta_{y_{m}}^{\prime}=\theta-\alpha\nabla_{\theta}\mathcal{L}_{y_{m}}(f_{\theta},\mathcal{D}^{spt}_{y_{m}}),$ (1)
    where $\mathcal{L}$ is the cross-entropy loss; $\alpha$ is the local learning rate; and $\theta_{y_{m}}^{\prime}$ is the locally updated recommender parameters on each task.
  • Global update: we calculate the testing loss on each $\mathcal{D}^{qry}_{y_{m}}$ with the corresponding $\theta_{y_{m}}^{\prime}$. The MAML global updating aims to update the initialization $\theta$ by one gradient step on the sum of all the testing losses:
    $\theta=\theta-\beta\nabla_{\theta}\sum\mathcal{L}_{y_{m}}(f_{\theta^{\prime}_{y_{m}}},\mathcal{D}^{qry}_{y_{m}}),$ (2)
    where $\beta$ is the global learning rate.
The optimal parameters of the meta-learning model can be obtained through alternating optimization of the inner and outer loops. Therefore, the goal of the meta-optimization is to minimize the loss function of the task:
$\min_{\theta}\sum\mathcal{L}_{y_{m}}(f_{\theta^{\prime}_{y_{m}}})=\sum\mathcal{L}_{y_{m}}(f_{\theta}-\alpha\nabla_{\theta}\mathcal{L}_{y_{m}}(f_{\theta})).$ (3)
II-A2 Related Task-aware Meta-learning
Directly transferring patterns from the auxiliary tasks to the target task may introduce noise thus adversely affecting the model performance. By holding the principle of “paying more attention to more relevant knowledge”, we further consider the relevance of different working conditions in different tasks when conducting the global update. We obtain the tasks relevance (e.g., $\gamma_{rel}$) based on different vibration signal patterns, we thus attentively adapt the gradient across tasks by employing their relevance.
To be specific, we employed autoencoders, a class of unsupervised neural networks, to ascertain the relevance among diverse auxiliary tasks with target task. For three distinct auxiliary tasks as example, represented by datasets $\mathcal{Y}_{0}$, $\mathcal{Y}_{1}$, and $\mathcal{Y}_{2}$, and a target task $\mathcal{Y}_{tar}$, an autoencoder was trained on their amalgamated data. This model comprises two parts: the encoder $f_{\text{enc}}$ and the decoder $f_{\text{dec}}$. In essence, the encoder compresses the input $x$ into a latent representation $y$, and the decoder attempts to reconstruct $x$ from $y$. Post-training, the encoder transforms each dataset into this compact latent space, yielding representations $\mu$. For a holistic view of each task, we calculate the mean vector of these representations, denoted as $\gamma_{rel}$. The relevance between auxiliary tasks and target task is then quantified using the Euclidean distance:
$\gamma_{rel_{i,t}}=\frac{1}{\sqrt{1+\sum_{k}(\mu_{i,k}-\mu_{t,k})^{2}}}$ (4)
where indices $i$ designate different auxiliary tasks, $t$ is the target task and $k$ traverses the latent dimensions. The derived distances, namely $\gamma_{rel_{0,t}}$, $\gamma_{rel_{1,t}}$, and $\gamma_{rel_{2,t}}$, serve as task relevance, with values closer to 1 suggesting more related tasks. In other words, if the auxiliary task is more related to the target task, we adapt the gradient so that it updates faster in that direction. There, the Eq.(1) is reformulated as:
$\theta_{y_{m}}^{\prime}=\theta-\alpha\nabla_{\theta}[\mathcal{L}_{y_{m}}(f_{\theta},\mathcal{D}^{spt}_{y_{m}})\times\gamma_{rel}].$ (5)
II-A3 Curriculum Strategy
In light of the challenges presented by high task diversity, we introduce a novel strategy inspired by the curriculum learning paradigm to enhance the convergence rate and generalization capacity of the meta-learner. We advocate for a sequential presentation of tasks to the meta-learner, adhering to an “easier first, harder later” doctrine. To operationalize this, we introduce a teacher recommender. For each auxiliary task, the recommender undergoes preliminary training and uses the highest validation score as an indicator of task difficulty. This metric subsequently informs the design of a curriculum for task sampling at each meta-training iteration. More specific, to quantify the difficulty of a given task, we partition the samples into a training set $D_{\text{train}}$ and a validation set $D_{\text{valid}}$. We train a base LSTM model extensively on $D_{\text{train}}$ as tearch recommender. Using the performance metric $\Phi$ (e.g., AUC) on the validation set, we ascertain the task-level difficulty, denoted as $\delta_{c}$. Specifically, $\delta_{c}$ is inversely related to the highest validation score achieved over all model parameters $\theta$:
$\delta_{c}\propto-\max_{\theta}\Phi(f_{\theta},D_{\text{valid}_{c}}|D_{\text{train}_{c}})$ (6)
Intuitively, a lower peak validation score indicates more challenging vibration signal pattern within a task.
II-B Target-level Fine-tune
After obtaining the general pattern of auxiliary tasks through the meta-learning paradigm, it is important to further fine-tune the learning model for the target task. By investigating the freezing of layers and then fine-tuning, the network generalizes better than one trained directly on the target dataset [24]. This operation not only incorporates information from the auxiliary tasks but also allows for better adaptation to the target task [21].
Specifically, assuming the model (e.g., LSTM) through meta-level training contains $L$ layers, we freeze first $l$ $(1\leq l\leq L)$ layers, while adding $n$ layers after the $l$ layers and fine tune the parameters by using data of the target task, so as to maximize the reusability of the general parameters. Hence, $\mathbf{e}_{hist}^{(tar)}$ is fed into the recurrent layer to infer the hidden state $\boldsymbol{h^{u}_{t_{k}}}$ at $t_{k}$, given by,
$\boldsymbol{h^{u}_{t_{k}}}=LSTM_{frozen}(\mathbf{e}_{hist}^{(tar)}).$ (7)
By doing this, the freezing-layer operation helps generate a network, capable of better balancing the parameters between the auxiliary tasks and target task after the final model fine-tuning.
The probability distribution on all working conditions is calculated by softmax function:
$\boldsymbol{\hat{y}}=softmax(f(\mathbf{h}^{u}_{t_{k}}))$ (8)
where $f$ is a fully connected layer to transform $\boldsymbol{h^{u}_{t_{k}}}$ into a $|\mathcal{P}|$-dimensional vector, and $|\mathcal{P}|$ is the number of faulty class in the target task. Hence, the objective function is defined by:
$\mathcal{J}=-\sum_{i=1}^{|\mathcal{P}|}\mathbf{y}[i]\cdot log(\hat{\mathbf{y}}[i])$ (9)
where $\mathbf{y}$ is an one-hot embedding of the ground-truth fault. Algorithm 1 summarizes the training process of CMEFD, consisting of meta training (lines 3-11), freezing layers and model fine tuning (lines 12-14), as well as fault diagnosis (lines 15-16).
Algorithm 1 Related Task Aware Curriculum Meta-learning (RT-ACM) Enhanced Fault Diagnosis framework
0:  $\mathbb{D}_{meta}^{(aux)}$; learning rates $\alpha$, $\beta$; number of shots $N$; max step of iterations $n$;
1:  Randomly initialize parameters $\theta$;
2:  Calculate the relevance between tasks by Eq.(4);
3:  Calculate the difficulty of auxiliary tasks by Eq.(6);
4:  while not done do
5:     for  $all$ $\mathbb{D}_{i}\in\mathbb{D}_{meta}^{(aux)}$ do
6:        Sample $N$ vibration signal from $\mathbb{D}_{i}$ as the adapt_batch based on the difficulty;
7:        Evaluate: $\nabla_{\theta}\mathcal{L}_{y_{m}}(f_{\theta},\mathcal{D}^{spt}_{y_{m}})$ using adapt_batch;
8:        Calculate the gradient update of $\theta^{\prime}_{y_{m}}$ by Eq.(5);
9:        Sample another $N$ vibration signal from $\mathbb{D}_{i}$ as the eval_batch based on the difficulty;
10:     end for
11:     Update $\theta$ using eval_batch by Eq.(2);
12:  end while
13:  Freeze the first $l$ layers as new LSTM model, i.e., LSTMfrozen;
14:  Fine-tune LSTMfrozen using only the target data $\mathbb{D}_{train}^{(tar)}$;
15:  Predict next possible fault via Eq.(8);
16:  Calculate the prediction loss for each record via Eq.(9);
III Case Study
TABLE I: COMPARISON OF TRADITIONAL DL-BASED AND META-LEARNING-BASED MODELS
(*The best results are highlighted in bold.)
3-way 6-way
Models 1-shot 5-shot 1-shot 5-shot
T1 T2 T3 T4 T1 T2 T3 T4 T1 T2 T3 T4 T1 T2 T3 T4
Deep Learning Based
LSTM 34.65 32.80 37.80 40.75 52.73 42.97 46.46 51.95 20.57 18.32 25.71 27.95 32.86 27.45 30.99 34.67
Cap-Net 41.23 35.56 45.24 47.89 69.53 55.43 65.31 63.74 37.53 31.83 41.02 44.65 63.18 50.29 59.55 57.38
WDCNN 41.50 34.48 45.90 46.07 69.84 55.75 66.93 63.23 37.82 30.92 42.45 44.03 63.68 50.53 60.29 57.05
Meta Learning Based
MAML 85.23 80.92 83.95 85.80 97.71 95.72 97.03 96.75 70.25 66.38 69.22 71.45 73.59 70.87 72.13 71.68
WCFS 96.77 95.52 97.03 97.70 98.70 97.59 98.88 98.73 85.67 83.32 89.15 87.49 89.68 87.97 93.27 90.15
CHAML 92.95 90.64 93.22 94.86 95.00 93.50 96.21 96.00 88.41 86.61 88.31 90.17 92.67 91.29 90.16 93.61
RT-ACM 98.51 97.44 98.62 98.53 99.68 99.53 99.73 99.56 93.22 92.49 94.38 95.14 97.25 96.59 97.48 97.05
Rolling bearings play a vital role as essential components in electric machines and find extensive usage across various mechanical equipment. Given the prolonged rotational wear and exposure to high temperatures during operation, bearing failures are common faults that occur. Therefore, conducting research on fault diagnosis in rolling bearings holds significant importance. This study specifically focuses on the detection of faults in rolling bearings within electric motor, considering the varying working conditions and a limited number of available samples. To evaluate the effectiveness of the proposed model, a series of case studies are conducted using two widely recognized datasets: the Case Western Reserve University (CWRU) Bearing Dataset [25], the Paderborn University Rolling Bearing Dataset [26] and the Gearbox dataset [19].
Moreover, several state-of-the-art models have been applied for comparison, including LSTM [23], the capsule network (Cap-Net) [4], wide deep CNN (WDCNN) [27], MAML [17], WDCNN-few-shot (WCFS) [10], CHAML [20], PKAML [19]. The LSTM, Cap-Net and WDCNN are DL-based model, MAML, WCFS, CHAML and PKAML are meta-learning based models.
(a) Healthy
(b) Ball 0.007
(c) Inner Race 0.007
(d) Outer Race 0.007
Figure 2: The visualisation of the vibration signals corresponding to different fault types under same fault diameter and working condition.
III-A Case 1: CWRU Bearing Dataset
III-A1 Dataset Description
The CWRU dataset has been commonly used to verify DL fault diagnosis methods. The selected rolling CWRU bearing dataset encompasses a comprehensive range of fault types, as depicted in Table II. The dataset comprises four distinct types of electric machine health states, which encompass one normal state and three fault states. The sizes of the faults are specified in terms of diameters, measuring 0.007, 0.014, and 0.021 inches. Fig. 2 provides a visual representation of the vibration signals corresponding to different fault types, all with the same fault diameter. As mentioned in Section II, fault diagnosis under various working conditions can be regarded as different tasks, and each task is a multi-class classification problem. Detailed descriptions of each task and relationship between the motor load and motor speed shown are presented in Table III. We treat one of the tasks as the target task and the rest as auxiliary tasks each time. Following [28], we chronologically divide the dataset of the target task into training, validation, and test sets with a ratio of 8:1:1.
TABLE II: INFORMATION OF USED BEARINGS FROM CWRU DATASET
Class Label Fault Type Fault Diameter (Inches)
1 Healthy 0
2 Ball 0.007
3 Ball 0.014
4 Ball 0.021
5 Inner Race 0.007
6 Inner Race 0.014
7 Inner Race 0.021
8 Outer Race 0.007
9 Outer Race 0.014
10 Outer Race 0.021
TABLE III: DETAILED DESCRIPTIONS OF EACH TASK UNDER CWRU DATASET
Task Load (hp) Number of Class Motor Speed (r/min)
$T^{1}$ 0 10 1797
$T^{2}$ 1 10 1772
$T^{3}$ 2 10 1750
$T^{4}$ 3 10 1730
III-A2 Hyper-parameter Settings
The optimal hyper-parameter settings for all models are empirically found out based on the performance on the validation set. Specifically, for TL-based moded, the learning rate is selected from {0.1, 0.05, 0.01, 0.005, 0.001, 0.0001}, and batch size is set from {256, 128, 64}. For meta-learning based models, the learning rate $\alpha,\beta$ are searched from {0.1, 0.01, 0.001, 0.0001}; and the batch size is set from {256, 128, 64} for a fair comparison. For RT-ACM, the number of freezing laysers $l$ is searched in range of [1,4] stepped by one, where the best setting is 3 for all tasks.
Figure 3: Accuracy in the (a) outer loop and (b) inner loop.
Figure 4: Model performance on $T^{1}$.
III-A3 Performance Comparison
Table I presents a comprehensive comparison of traditional deep learning-based models and meta-learning-based models across various tasks and shots. The results indicate a clear performance superiority of meta-learning-based models over their traditional counterparts. Specifically, our method RT-ACM consistently achieves the highest accuracy across all scenarios, notably outperforming other models with an impressive accuracy range of 98.51% to 99.73% in the 3-way tasks and 93.22% to 97.48% in the 6-way tasks. MAML, WCFS and CHAML, while also meta-learning-based, demonstrate strong performances, especially in the 3-way 5-shot settings. CHAML generally outperforms WCFS due to its structured learning process. By leveraging a curriculum sampling, CHAML progressively exposes the model to increasingly complex tasks, enhancing its ability to generalize and perform well across diverse scenarios. This staged learning process allows the model to build a more robust understanding of the task at hand, resulting in higher accuracy, particularly in more challenging 6-way tasks. Our proposed method RT-ACM integrates related task-aware meta-learning with a curriculum strategy, which further amplifies its performance. The related task-aware mechanism ensures that the model is adept at identifying and leveraging similarities between tasks, thereby improving its efficiency in learning new tasks. Additionally, the curriculum strategy within RT-ACM helps in systematically increasing the difficulty of the tasks presented during training, allowing the model to develop a deeper and more nuanced understanding of the task space. This dual approach not only enhances the learning process but also ensures that the model remains adaptable and highly accurate, as evidenced by its superior performance across all evaluated metrics.
Fig. 3a shows the performance comparison of four meta-learning-based models, in terms of accuracy over increasing global update steps. Our proposed method, RT-ACM, which leverages related task-aware meta-learning and a curriculum strategy, demonstrates a superior ability to rapidly achieve high accuracy. This rapid improvement can be attributed to the efficient transfer of knowledge from related tasks and a structured learning process that progresses from simpler to more complex tasks. In contrast, CHAML, which employs curriculum sampling, lags behind RT-ACM and fails to match its performance throughout the training process due to a less effective related task prioritization. WCFS initially shows competitive accuracy but experiences a decline in performance in later stages, possibly due to overfitting or poor generalization. MAML, while maintaining steady progress, does not achieve the same level of accuracy as RT-ACM. Fig. 3b illustrates the relationship between the performance of model and the number of local update steps. Through empirical analysis, it is observed that a single-step update is adequate to enhance the accuracy significantly.
Fig. 4 illustrates that the RT-ACM model stands out for its exceptional combination of efficiency and accuracy on 3-way 5-shot. When compared to conventional models like LSTM and WDCNN, RT-ACM’s testing time is considerably shorter,which reflecting its higher processing speed. Additionally, RT-ACM competes closely in accuracy with more sophisticated models like MAML and WCFS. These results highlight the potential of RT-ACM, underscoring the advantages of incorporating auxiliary tasks to enhance the target task in terms of accuracy. Moreover, the curriculum transfer of knowledge from auxiliary tasks in RT-ACM demonstrates rapid learning capabilities.
(a) LSTM
(b) MAML
(c) WCFS
(d) RT-ACM
Figure 5: The visualization of classification from four models.
(a) LSTM
(b) MAML
(c) WCFS
(d) RT-ACM
Figure 6: The visualization of confusion matrices from four models.
III-B Case 2: PU Bearing Dataset
III-B1 Dataset Description
Different from CWRU dataset, the Paderborn University (PU) Bearing Dataset is a pivotal diagnostic tool for rolling bearings include the artificially damaged data. To mitigate the influence of similar bearings on the dataset, vibration signals were collected from 26 bearings characterized by artificial damage, natural damage, and healthy states. The detailed working conditions are presented in Table IV. Following [19], the first 20 bearings were utilized for training, while the remaining 6 bearings were allocated for testing. Each bearing exhibited four distinct working conditions are presented in Table V, primarily varying in rotational speed, load torque, and radial force.
TABLE IV: INFORMATION OF USED BEARINGS FROM PU DATASET
No. Bearing Code Faulty Type Damage Method
1 K003 Healthy -
2 K004 Helathy -
3 K005 Healthy -
4 K006 Healthy -
5 KA01 Outer Race Muliple Damage
6 KA03 Outer Race Electric Engraver
7 KA05 Outer Race Electric Engraver
8 KA07 Outer Race Drilling
9 KA08 Outer Race Drilling
10 KI01 Inner Race Muliple Damage
11 KI05 Inner Race Electric Engraver
12 KI07 Inner Race Electric Engraver
13 KA15 Outer Race Plastic Deform
14 KA16 Outer Race Fatigue
15 KA22 Outer Race Fatigue
16 KB23 Inner/Outer Fatigue
17 KB27 Inner/Outer Plastic Deform
18 KI16 Inner Race Fatigue
19 KI17 Inner Race Fatigue
20 KI21 Inner Race Fatigue
21 K001 Healthy -
22 K002 Healthy -
23 KA30 Outer Race Plastic Deform
24 KB24 Inner/Outer Fatigue
25 KI14 Inner Race Fatigue
26 KI18 Inner Race Fatigue
TABLE V: DETAILED DESCRIPTIONS OF EACH TASK UNDER PU DATASET
Task Rational speed (rpm) Load torque (Nm) Radial force (N)
$T^{1}$ 1500 0.7 400
$T^{2}$ 900 0.7 1000
$T^{3}$ 1500 0.1 1000
$T^{4}$ 1500 0.7 1000
TABLE VI: THE COMPARATIVE METHODS IN PU DATASET ($T^{1}$)
(*The best results are highlighted in bold.)
Models Accuracy Precision F1-score
10-way
1-shot 5-shot 1-shot 5-shot 1-shot 5-shot
LSTM 40.15 50.89 39.92 50.05 39.61 49.84
MAML 76.85 85.35 76.71 85.19 76.63 85.12
WCFS 78.52 91.07 78.37 90.92 78.45 91.03
CHAML 78.75 90.20 77.61 91.07 78.58 91.13
PKAML 79.20 92.09 79.13 91.97 79.16 92.01
RT-ACM 81.57 94.34 81.43 94.21 81.46 94.27
(a) Training Loss
(b) Validation Accuracy
Figure 7: The visualization of training loss and validation accuracy from four models.
III-B2 Evaluation Metrics
Given the inclusion of artificially damaged data in the PU dataset, resulting in imbalanced distribution, precision and F1 score offer better insights into model performance. Thus, three established evaluation metrics are employed [6]: Accuracy, F1-Score and Precision. Accuracy quantifies the proportion of correct predictions in the total predictions made. It is given by the formula:
Accuracy=Number of Correct PredictionsTotal Number of $\text{Accuracy}=\frac{\text{Number of Correct Predictions}}{\text{Total Numberof Predictions}}$
Accuracy provides a general measure of model performance, especially when the class distributions are balanced. Precision is the fraction of relevant instances among the retrieved instances. In other words, it quantifies the accuracy of positive predictions. It is given by:
Precision=True PositivesTrue Positives+False $\text{Precision}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Positives}}$
Precision provides insight into the reliability of positive predictions made by the model. The F1 score is the harmonic mean of precision and recall. It balances the trade-off between precision and recall and is especially useful when the class distribution is imbalanced. It is computed as:
[…]
Where, precision is the proportion of true positive predictions in the set of positive predictions made. Recall is the proportion of true positive predictions in the set of actual positives.
III-B3 Performance Comparison
Table VI shows a performance comparison of various models on the PU dataset in 10-way, 1-shot and 5-shot. Notably, all methods except LSTM are meta-learning based approaches, allowing them to handle data sparsity issues more effectively than LSTM.. Among these methods, PKAML employs a pretraining process to integrate prior knowledge, enhancing its performance. CHAML utilizes a curriculum sampling in meta-training, providing a structured approach to learning. However, both PKAML and CHAML do not adequately consider the diversity and relevance of auxiliary tasks in relation to the target task, which can limit their effectiveness. In contrast, our proposed method, RT-ACM, incorporates related task-aware meta-learning along with a curriculum strategy, resulting in superior performance. This is evidenced by RT-ACM achieving the highest scores across all metrics and shots, with an accuracy of 81.57% and 94.34% in 1-shot and 5-shot settings, respectively, outperforming the runner-up PKAML and other models significantly.
Fig. 5 presents t-SNE visualizations for four models, highlighting their ability to discriminate between classes. The LSTM visualization shows poorly separated clusters, indicating significant class overlap and lower classification efficacy. MAML improves with tighter clusters but still suffers from overlap at the cluster centers. WCFS slightly betters class separation, with clearer clusters but some outliers. RT-ACM’s visualization stands out, with each class forming distinct, isolated clusters, showcasing superior class discrimination and robust classification. The color-coded clusters denote various classes, and RT-ACM’s close-knit clusters suggest strong internal similarity and external dissimilarity among classes, reflecting RT-ACM’s effective feature space transformation for enhanced class distinction.
Fig. 6 showcases confusion matrices for four classification model, illustrating each model’s predictive performance. The LSTM matrix shows widespread misclassifications, such as confusing class 2 with class 3, indicating difficulty in distinguishing similar classes. MAML improves, concentrating predictions along the diagonal but still mixes up classes like 4 and 5. WCFS further enhances accuracy, with clearer diagonal values and fewer errors, notably between classes 7 and 8. RT-ACM stands out with a strong diagonal focus, demonstrating superior accuracy and minimal misclassification, especially distinguishing between closely related classes like 5 and 6, and 9 and 10, where other models struggle. This highlights RT-ACM’s robust predictive capability across diverse classes.
Fig. 7a illustrates the training loss of various models, including LSTM, MAML, WCFS, and RT-ACM. The LSTM exhibits significant fluctuations, indicating instability from a small dataset size and potential overfitting issues. In contrast, RT-ACM shows a quick and steady reduction in loss, outperforming MAML and WCFS in terms of learning efficiency and model regularization, suggesting a superior dataset fit. Fig. 7b compares validation accuracy, with RT-ACM achieving higher accuracy early and maintaining it, highlighting its robustness and generalization capability. LSTM’s accuracy fluctuates, while MAML and WCFS improve steadily but don’t match RT-ACM’s performance. RT-ACM excels in convergence speed and generalization, benefiting from task-aware meta-learning and a curriculum strategy.
IV Conclusion and Future Work
In this paper, we develop the Related Task Aware Curriculum Meta-learning (RT-ACM) enhanced fault diagnosis framework, a novel approach in the realm of fault diagnosis for smart manufacturing designed to address the issue of data scarcity. RT-ACM innovatively utilizes auxiliary tasks to enrich the primary fault diagnosis tasks, adhering to the principle of “paying more attention to more relevant knowledge” by focusing on the most relevant knowledge and adopting an “easy-to-hard” curriculum learning strategy. Our comprehensive evaluations across three real-world datasets confirm the effectiveness and superiority of the RT-ACM framework. The experiments not only highlight the efficacy of the framework but also validate the architectural choices, underscoring the potential of combining curriculum strategy with related task aware meta-learning for enhanced fault diagnosis. In future work, we will explore new strategies that can learn both temporal and spatial representations of auxiliary tasks when transferring knowledge to the target task.

### v2

@@
-Leveraging Auxiliary Task Relevance for Enhanced Industrial Fault Diagnosis through Curriculum Meta-learning
+Leveraging Auxiliary Task Relevance for Enhanced Bearing Fault Diagnosis through Curriculum Meta-learning
 Abstract
@@
 To address these challenges, a Related Task Aware Curriculum Meta-learning (RT-ACM) enhanced fault diagnosis framework is proposed in this paper, inspired by human cognitive learning processes.
-RT-ACM improves training by considering the relevance of auxiliary working conditions, adhering to the principle of “paying more attention to more relevant knowledge”, and focusing on “easier first, harder later” curriculum sampling.
+RT-ACM improves training by considering the relevance of auxiliary sensor working conditions, adhering to the principle of “paying more attention to more relevant knowledge”, and focusing on “easier first, harder later” curriculum sampling.
 This approach aids the meta-learner in achieving a superior convergence state.
@@
 I Introduction
-Fault diagnosis is crucial for ensuring the safe and efficient operation of machines [1].
-Recently, methods based on deep learning (DL) have demonstrated considerable promise in accurately diagnosing machine faults [2].
+Fault diagnosis is crucial for ensuring the safe and efficient operation of machines [1, 2].
+Recently, methods based on deep learning (DL) have demonstrated considerable promise in accurately diagnosing machine faults through sensor signals [3, 4].
 An advanced ensemble convolutional neural network (CNN) has been introduced for this purpose, utilizing diverse local minima to enhance model adaptability [5].
@@
 Three key strategies have emerged to tackle the challenge of data scarcity in creating reliable fault diagnosis models: Data Augmentation (DA) [10], Transfer Learning (TL) [11], and Few-Shot Learning (FSL) approaches [12].
-At its essence, DA aims to expand and diversify datasets by producing additional data from the limited original samples available.
-For example, data enhancement techniques that incorporate Gaussian noise, masking noise, and signal translation have been explored in [8].
-Moreover, generative models, including variational autoencoders and generative adversarial networks, have been employed to generate fault data for training DL-based diagnostic models in [11].
+DA aims to expand and diversify datasets by producing additional data from the limited original samples available.
+For example, data enhancement techniques that incorporate Gaussian noise, masking noise, and signal translation have been explored in [10, 13].
+Moreover, generative models, including variational autoencoders and generative adversarial networks, have been employed to generate sensor fault data for training DL-based diagnostic models in [14].
 However, a notable limitation of DA is the potential degradation in data quality and precision, especially when original datasets are sparse.
@@
 To circumvent the limitations of DA, the use of prior knowledge from related tasks has been explored, leading to the concept of Transfer Learning (TL).
-TL methods strive to utilize knowledge derived from one task to benefit a different, albeit related, task.
-For instance, a comprehensive one-dimensional CNN architecture, supplemented with fine-tuning mechanisms specifically for fault diagnosis, has been presented[1].
+TL methods aim to transfer knowledge gained from one task to improve performance on another closely related task.
+For instance, a comprehensive one-dimensional CNN architecture, supplemented with fine-tuning mechanisms specifically for fault diagnosis, has been presented[1, 15].
 Similarly, a CNN-based model for diagnosing faults in bearings has been proposed, with its basis being knowledge garnered from laboratory bearings data[16].
@@
 FSL delves deeper by leveraging insights from associated tasks to benefit the primary FSL task [19].
-In practical settings, machines typically operate under varied conditions.
-Gathering substantial data for each specific condition is both expensive and, in some cases, impractical.
+In practical settings, machines typically operate under varied conditions, gathering substantial sensor data for each specific condition is both expensive and, in some cases, impractical [20].
 Yet, when data from varied operational conditions are utilized, a multitude of auxiliary tasks can be formulated.
 This inherent flexibility renders meta-learning especially fitting and potent for few-shot fault diagnosis scenarios.
-Such FSL models specifically designed for fault diagnosis have been proposed in [16].
-These models frequently incorporate metric-based meta-learning methods, with Siamese neural networks and matching networks being prominent examples.
-Despite the potential of metric-based meta-learning methods for feature extraction, their efficacy is hindered by data sparsity and variations in auxiliary task distributions.
-Model-Agnostic Meta-Learning (MAML) [17], a optimization-based meta-learning method, diverges from traditional metric-based approaches by emphasizing the development of models with robust generalization capabilities.
-MAML aims to find an initial parameters set, then the model could achieve optimal performance on a series of new tasks after one or more gradient updates, which are calculated using a small amount of data from the series of tasks.
-Through minimal fine-tuning with target task data, MAML demonstrates superior classification performance in few-shot bearing fault diagnosis.
-Zhang et al.
-[18] proposed a MAML-based few-shot learning framework for bearing fault diagnosis with limited data.
-They achieved this by developing a fault diagnosis model using an improved meta-relation network.
+Such FSL models specifically designed for fault diagnosis have been proposed in [21, 22].
+These models frequently incorporate metric-based meta-learning methods.
+Metric-based meta-learning methods have demonstrated potential in feature extraction; however, their effectiveness is often limited by data sparsity and variations in auxiliary task distributions.
+Model-Agnostic Meta-Learning (MAML) [23], an optimization-based meta-learning framework, offers a different approach by focusing on developing models with strong generalization capabilities.
+MAML seeks to identify a set of initial parameters that can be fine-tuned with minimal data to achieve optimal performance across a series of new tasks.
+As shown in Fig.
+1, MAML optimizes the model to establish a representation, $\theta$, that facilitates rapid adaptation to new tasks.
+By leveraging a small amount of target task data for fine-tuning, MAML achieves notable classification accuracy in few-shot bearing fault diagnosis.
+For instance, Zhang et al.
+[24] proposed a MAML-based framework employing an enhanced meta-relation network for bearing fault diagnosis with limited data.
 Zhou et al.
-[19] introduced an augmented method that integrates prior knowledge into MAML to fully utilize the prior knowledge present in both training and testing sets.
+[25] further advanced this concept by integrating prior knowledge into MAML, enabling more effective utilization of prior knowledge in both training and testing phases.
 Chen et al.
-[10] proposed a MAML based model featuring a four-layer CNN capable of feature extraction and rapid adaptation, enabling it to quickly adapt to fault diagnosis tasks under various working conditions.
+[12] developed a MAML-based model featuring a four-layer CNN designed for feature extraction and rapid adaptation, allowing efficient fault diagnosis under diverse working conditions.
+Additionally, Cui et al.
+[26] employed a federated meta-learning approach combining representation encoding with MAML to enhance performance in distributed fault diagnosis tasks.
 Although these studies have achieved some performance improvements and emphasized that MAML learns to learn, they have overlooked the difficulty level of each task and have not considered the human learning experience of easy-to-hard curriculum learning.
-Chen et al.
-[20] propose a novel curriculum hardness aware meta-Learning framework, which incorporates hard sample mining and curriculum learning into a meta-learning paradigm.
-Hu et al.
-[6] utilized a specialized task-sequencing strategy for few-shot fault diagnosis, introducing a method where auxiliary tasks are arranged from simple to complex to improve knowledge adaptability during meta-learning.
-Nevertheless, these existing studies have not considered the relevance between different tasks and the target task.
-Irrelevant auxiliary tasks may negatively impact the target task.
-To simultaneously consider the human easy-to-hard curriculum learning processing and the relevance between different tasks and the target task, we propose a novel Related Task Aware Curriculum Meta-learning (RT-ACM) enhanced fault diagnosis framework, which delicately considers the relevance of patterns among different tasks within the meta-learning paradigm to alleviate data sparsity and task diversity issues.
-To begin with, RT-ACM extends MAML to facilitate knowledge transfer to the target task under conditions of limited data.
-With fault diagnosis in different working conditions treated as single tasks, the initial model weights are meta-learned for faster adaptation to limited training data in target tasks.
-Furthermore, we design two components in the meta-training process: related task-aware meta-learning and a curriculum strategy to explicitly address the task diversity issue.
-Firstly, we exploit the heuristic idea of “pay more attention to more relevant knowledge” to enhance the meta-learner by considering the relevance of auxiliary tasks as related task-aware meta-learning [21].
-This approach compels the model to learn more from a combination of tasks that are more similar and are expected to be highly diverse, thereby contributing more to generalization.
-Specifically, we divide each meta-training round into two stages.
-In each stage, we first update the meta-learner in a MAML-like manner and then conditionally re-sample a new batch of tasks, focusing on the most complex working conditions based on validation score.
-Secondly, we propose drawing upon a curriculum strategy to improve the convergence rate and generalization capacity of the meta-learner given the high diversity among different tasks.
-The underlying concept is to present training tasks to the meta-learner in an “easier first, harder later” paradigm as determined by a teacher recommender [20, 22].
-This aims to guide the meta-learner towards a better state.
-Specifically, we pre-train a teacher recommender for each auxiliary task, use the best validation score to gauge task difficulty, and utilize this knowledge to design a curriculum for the task sampling pool at each meta-training step.
-Overall, our main contributions lie in three folds:
+A specialized task-sequencing strategy, task-sequencing meta-learning (TSML) [8], has been introduced for few-shot fault diagnosis.
+This approach arranges auxiliary tasks from simple to complex, aiming to enhance knowledge adaptability during meta-learning.
+However, it relies on k-means clustering to evaluate task difficulty, which overlooks the intricate relationships within temporal data, limiting its effectiveness to simpler datasets.
+Furthermore, these existing studies have not considered the relevance between different tasks and the target task, as unrelated auxiliary tasks can adversely affect the target task.
+Figure 1: Learning process of MAML, which optimizes for a representation $\theta$ that can quickly adapt to new tasks.
+To simultaneously consider data sparsity and task diversity challenges in fault diagnosis, we propose a novel Related Task Aware Curriculum Meta-learning (RT-ACM) framework that integrates human-inspired easy-to-hard curriculum learning with task relevance considerations.
+RT-ACM leverages the meta-learning paradigm to capture the relationships among patterns from diverse tasks, enabling more effective knowledge transfer under limited data conditions.
+Building on the foundation of MAML, the framework is designed to enhance adaptation to target tasks with minimal training data by treating fault diagnosis tasks under varying working conditions as individual tasks.
+The meta-training process incorporates two key components: a related task-aware meta-learning mechanism that emphasizes task relevance, and a curriculum strategy that systematically addresses task diversity, facilitating improved learning efficiency and generalization.
+To enhance the meta-learner, we exploit the heuristic idea of ”paying more attention to more relevant knowledge,” adopting a related task-aware meta-learning approach.
+This strategy encourages the model to prioritize learning from tasks that are highly relevant, thereby improving generalization.
+Each meta-training round is structured into two stages: first, the meta-learner is updated using a MAML-like process; second, a new batch of tasks is conditionally re-sampled, focusing on the most challenging working conditions as determined by validation scores.
+Specifically, we propose a curriculum strategy to refine the task sampling process during MAML training.
+Guided by the principle of an ”easier-first, harder-later” paradigm, this approach directs the meta-learner toward an optimized state.
+A teacher recommender is pre-trained for each auxiliary task, leveraging the best validation score to assess task difficulty and dynamically designing the task sampling curriculum at each meta-training step.
+Overall, our main contributions can be summarized in three key aspects:
   1. 1.
-    To the best of our knowledge, it is one of the early studies that examines to what extent knowledge can be transferred from auxiliary tasks to the target task by differentiating the relevance of working conditions.
+    To the best of our knowledge, this is one of the early studies that examines the extent to which knowledge can be transferred from auxiliary tasks to the target task by differentiating the relevance of working conditions.
+A novel related task-aware curriculum meta-learning (RT-ACM) enhanced fault diagnosis framework is proposed, which leverages both transferred knowledge and the diversity of tasks within the target task to mitigate the issue of data limitation.
   2. 2.
-    A novel Related Task Aware Curriculum Meta-learning (RT-ACM) enhanced fault diagnosis framework is proposed, which leverages both transferred knowledge and the diversity of tasks within the target task to mitigate the issue of data sparsity.
+    A meta-training process is formulated, which incorporates related task-aware meta-learning and a curriculum strategy to leverage auxiliary tasks while balancing the high diversity among different tasks.
+The related task-aware meta-learning follows the principle of “paying more attention to more relevant tasks”, while the curriculum strategy follows the principle of “easy first, hard later.” Moreover, we design a teacher recommender to evaluate task difficulty in temporal data.
   3. 3.
     Extensive experiments have been conducted on two real-world datasets to validate the superiority of RT-ACM against state-of-the-art approaches.
-The rest of this article is organized as follows.
-The proposed method is detailed in Section II.
-In Section III, the effectiveness of RT-ACM is evaluated with various case studies.
-Finally, Section IV concludes this article and presents future work.
+Furthermore, we have made our source code publicly available to contribute further to advancements in this field.
+The remainder of this article is structured as follows: Section II provides a detailed description of the proposed method.
+Section III presents a series of case studies to evaluate the effectiveness of RT-ACM.
+Finally, Section IV offers conclusions and outlines directions for future work.
 II Proposed RT-ACM
-Assume that $A$ auxiliary tasks are available and denoted by $\left\{T^{\mu}\right\}_{\mu=1}^{A}$.
-In each auxiliary task, $T^{\mu}=\left\{(x^{\mu}_{i},y^{\mu}_{i})\right\}^{M}_{i=1}$, where $x^{\mu}_{i}\in\mathcal{R}^{1\times D}$ is $i$th sample of the input data.
-In the field of bearing fault diagnosis, 1-D vibration signal is commonly used as the input data.
-The vibration signal is segmented using a sliding window to generate samples, and $D$ is the length of the sliding window.
-$y_{i}^{\mu}$ is the health status of $x_{i}^{\mu}$.
-$M$ denotes the number of samples contained in $T^{\mu}$.
-A target task is denote by $T^{tar}=\left\{(x^{tar}_{i},y^{tar}_{i})\right\}^{M^{tar}}_{i=1}$, where only a few samples are available for training.
-Thus, $M^{tar}$, the number of samples contained in $T^{tar}$, is assigned a very small value.
-$x^{tar}_{i}\in\mathcal{R}^{1\times D}$ is the input data of the $i$th sample and $y^{tar}_{i}$ is the corresponding label.
-The overview of RT-ACM is outlined in Fig.
-1, mainly composed of two channels (i.e., meta-level training on auxiliary tasks and target-level fine-tune).
-In particular, the meta-level training exploits meta-learning to capture the common vibration pattern by holding the principle of “paying more attention to more relevant knowledge” and learning knowledge curriculum judge by teacher recommender.
-The goal of target-level fine-tune is to learn the accurate vibration pattern in the target task and performs the predictions.
+Let $A$ auxiliary tasks be represented as $\left\{T^{\mu}\right\}_{\mu=1}^{A}$, where each auxiliary task $T^{\mu}$ comprises a dataset $\left\{(x^{\mu}_{i},y^{\mu}_{i})\right\}_{i=1}^{M}$.
+Here, $x^{\mu}_{i}\in\mathcal{R}^{1\times D}$ corresponds to the $i$th input sample derived from 1-D vibration sensor signals, a common data source in this domain.
+These signals are segmented using a sliding window of length $D$ to produce samples, and $y^{\mu}_{i}$ denotes the sensor health status label associated with $x^{\mu}_{i}$.
+The total number of samples in $T^{\mu}$ is given by $M$.
+Similarly, a target task is defined as $T^{tar}=\left\{(x^{tar}_{i},y^{tar}_{i})\right\}_{i=1}^{M^{tar}}$, where the number of available training samples, $M^{tar}$, is limited, reflecting the scarcity of labeled data.
+Each $x^{tar}_{i}\in\mathcal{R}^{1\times D}$ represents the input data for the $i$th sample, while $y^{tar}_{i}$ is its corresponding label.
+The structure of RT-ACM is depicted in Fig.
+2, consisting primarily of two components: meta-level training on auxiliary tasks and fine-tuning on the target task.
+Specifically, the meta-level training employs meta-learning to extract shared sensor vibration patterns, adhering to the principles of “paying more attention on more relevant knowledge” and “easy first, hard later,” as determined by the teacher recommender.
+Meanwhile, the target-level fine-tuning aims to capture precise vibration patterns within the target task to enhance prediction accuracy.
 Figure 2: The overall framework of our proposed RT-ACM.
 II-A Meta-level Training on Auxiliary Tasks
-To distil knowledge from auxiliary tasks and employ vibration patterns, we extend the MAML with long short-term memory (LSTM) [23] as the framework for the meta-learning update.
-In particular, we devise a related task aware meta-learning and curriculum strategy that can transfer knowledge based on the relevance of patterns among tasks.
+To distill knowledge from auxiliary tasks and leverage vibration patterns, we enhance the MAML framework by incorporating long short-term memory (LSTM) [27] for meta-learning updates.
+Notably, we propose a related task-aware meta-learning approach combined with a curriculum strategy, enabling knowledge transfer by evaluating the relevance of patterns across tasks.
 II-A1 Basic Theory of MAML
-Unlike traditional deep learning, as a category of meta-learning, the MAML algorithm is designed to train the initialization parameters of a model within the outer loop.
-By learning a set of initialization parameters that are sufficiently general, the model can achieve robust performance efficiently with only a few gradient update steps when adapting to new tasks with limited data.
-From a learning perspective, MAML seeks to identify general internal representations that facilitate transferability to various tasks.
-Consequently, when confronted with new tasks, only minimal parameter fine-tuning is necessary.
-Following [20], the fault diagnosis in different working conditions can be viewed as a single task in a meta-learning paradigm.
-Thus, the vibration signal of auxiliary tasks $\left\{T^{\mu}\right\}_{\mu=1}^{A}$ are denoted as $\mathbb{D}_{meta}^{(aux)}$ and the vibration signal of target tasks $T^{tar}$ are divided as training sets $\mathbb{D}_{train}^{(tar)}$ and test sets $\mathbb{D}_{test}^{(tar)}$.
-We treat each working conditions $y_{m}$ as a meta-learning task, where each task has support set $\mathcal{D}^{spt}_{y_{m}}$ for training and a query set $\mathcal{D}^{qry}_{y_{m}}$ for testing.
-Finally, our goal is to leverage the auxiliary tasks to learn a meta-learner $F_{w}$, which contains knowledge about different working conditions.
-Specifically, each iteration of MAML includes local update and global update on the sampled task batch, where the first phase updates $\theta$ locally on the $\mathcal{D}^{spt}$ of each task, and the second phase globally updates $\theta$ by gradient descent to minimize the sum of loss on the $\mathcal{D}^{qry}$ of all tasks.
-  • Local update: we first sample a batch of tasks from $\mathbb{D}^{(aux)}_{meta}$.
-Then we randomly sample a group of vibration signal from $\mathcal{D}^{spt}_{y_{m}}$ and $\mathcal{D}^{qry}_{y_{m}}$ for each sampled tasks.
-Thus, we calculate the training loss on $\mathcal{D}^{spt}_{y_{m}}$ and locally update $\theta$ by one step:
+As a subset of meta-learning, the MAML algorithm differs from traditional deep learning by focusing on training a model’s initialization parameters within its outer loop.
+By optimizing these initialization parameters to be broadly applicable, the model can adapt to new tasks with limited data efficiently, requiring only a few gradient updates.
+From a learning perspective, MAML aims to uncover general internal representations that promote effective transferability across diverse tasks.
+As a result, minimal parameter adjustments are needed when tackling new tasks.
+Following [17], fault diagnosis across various working conditions can be framed as a single task within the meta-learning framework.
+Accordingly, the vibration signals from auxiliary tasks $\{T^{\mu}\}{\mu=1}^{A}$ are represented as $\mathbb{D}_{meta}^{(aux)}$, while the vibration signals from target tasks $T^{tar}$ are partitioned into training sets $\mathbb{D}_{train}^{(tar)}$ and testing sets $\mathbb{D}_{test}^{(tar)}$.
+Each working condition $y_{m}$ is treated as an individual meta-learning task, with a support set $\mathcal{D}^{spt}_{y_{m}}$ for training and a query set $\mathcal{D}^{qry}_{y_{m}}$ for evaluation.
+Ultimately, the objective is to utilize the auxiliary tasks to train a meta-learner $F_{w}$ that encapsulates knowledge across various working conditions.
+Specifically, in each iteration of MAML, the process involves a local update and a global update on a batch of sampled tasks.
+During the local update phase, the model parameters $\theta$ are adjusted on the support set $\mathcal{D}^{spt}$ for each task.
+Subsequently, in the global update phase, $\theta$ is updated using gradient descent to minimize the aggregate loss computed on the query set $\mathcal{D}^{qry}$ across all tasks.
+  • Local Update: A batch of tasks is first sampled from $\mathbb{D}^{(aux)}_{meta}$.
+For each selected task, a subset of vibration signals is randomly drawn from its support set $\mathcal{D}^{spt}_{y_{m}}$ and query set $\mathcal{D}^{qry}_{y_{m}}$.
+The training loss is then computed on $\mathcal{D}^{spt}_{y_{m}}$, and the parameters $\theta$ are locally updated with a single gradient descent step:
     $\theta_{y_{m}}^{\prime}=\theta-\alpha\nabla_{\theta}\mathcal{L}_{y_{m}}(f_{\theta},\mathcal{D}^{spt}_{y_{m}}),$ (1)
     where $\mathcal{L}$ is the cross-entropy loss; $\alpha$ is the local learning rate; and $\theta_{y_{m}}^{\prime}$ is the locally updated recommender parameters on each task.
-  • Global update: we calculate the testing loss on each $\mathcal{D}^{qry}_{y_{m}}$ with the corresponding $\theta_{y_{m}}^{\prime}$.
-The MAML global updating aims to update the initialization $\theta$ by one gradient step on the sum of all the testing losses:
+  • Global Update: The testing loss is evaluated on each $\mathcal{D}^{qry}_{y_{m}}$ using the corresponding locally updated parameters $\theta_{y_{m}}^{\prime}$.
+The goal of the MAML global update is to update the initialization $\theta$ by performing a single gradient descent step on the aggregated testing losses across all tasks:
     $\theta=\theta-\beta\nabla_{\theta}\sum\mathcal{L}_{y_{m}}(f_{\theta^{\prime}_{y_{m}}},\mathcal{D}^{qry}_{y_{m}}),$ (2)
     where $\beta$ is the global learning rate.
-The optimal parameters of the meta-learning model can be obtained through alternating optimization of the inner and outer loops.
-Therefore, the goal of the meta-optimization is to minimize the loss function of the task:
+The optimal parameters for the meta-learning model are achieved through iterative optimization within the inner (local update) and outer (global update) loops.
+Consequently, the objective of meta-optimization is to minimize the task-specific loss function:
 $\min_{\theta}\sum\mathcal{L}_{y_{m}}(f_{\theta^{\prime}_{y_{m}}})=\sum\mathcal{L}_{y_{m}}(f_{\theta}-\alpha\nabla_{\theta}\mathcal{L}_{y_{m}}(f_{\theta})).$ (3)
 II-A2 Related Task-aware Meta-learning
-Directly transferring patterns from the auxiliary tasks to the target task may introduce noise thus adversely affecting the model performance.
-By holding the principle of “paying more attention to more relevant knowledge”, we further consider the relevance of different working conditions in different tasks when conducting the global update.
-We obtain the tasks relevance (e.g., $\gamma_{rel}$) based on different vibration signal patterns, we thus attentively adapt the gradient across tasks by employing their relevance.
+Directly transferring patterns from auxiliary tasks to the target task can introduce noise, potentially degrading model performance.
+To address this, we adhere to the principle of “paying more attention on more relevant knowledge” and incorporate the relevance of working conditions across tasks during the global update.
+By calculating task relevance (e.g., $\gamma_{rel}$) based on variations in vibration signal patterns, we adapt gradients across tasks with attention to their relevance.
 To be specific, we employed autoencoders, a class of unsupervised neural networks, to ascertain the relevance among diverse auxiliary tasks with target task.
@@
 The derived distances, namely $\gamma_{rel_{0,t}}$, $\gamma_{rel_{1,t}}$, and $\gamma_{rel_{2,t}}$, serve as task relevance, with values closer to 1 suggesting more related tasks.
-In other words, if the auxiliary task is more related to the target task, we adapt the gradient so that it updates faster in that direction.
-There, the Eq.(1) is reformulated as:
+In essence, if an auxiliary task exhibits greater relevance to the target task, the gradient is adjusted to facilitate faster updates in that direction.
+Consequently, Eq.
+(1) is revised as follows:
 $\theta_{y_{m}}^{\prime}=\theta-\alpha\nabla_{\theta}[\mathcal{L}_{y_{m}}(f_{\theta},\mathcal{D}^{spt}_{y_{m}})\times\gamma_{rel}].$ (5)
 II-A3 Curriculum Strategy
-In light of the challenges presented by high task diversity, we introduce a novel strategy inspired by the curriculum learning paradigm to enhance the convergence rate and generalization capacity of the meta-learner.
-We advocate for a sequential presentation of tasks to the meta-learner, adhering to an “easier first, harder later” doctrine.
-To operationalize this, we introduce a teacher recommender.
-For each auxiliary task, the recommender undergoes preliminary training and uses the highest validation score as an indicator of task difficulty.
-This metric subsequently informs the design of a curriculum for task sampling at each meta-training iteration.
-More specific, to quantify the difficulty of a given task, we partition the samples into a training set $D_{\text{train}}$ and a validation set $D_{\text{valid}}$.
-We train a base LSTM model extensively on $D_{\text{train}}$ as tearch recommender.
-Using the performance metric $\Phi$ (e.g., AUC) on the validation set, we ascertain the task-level difficulty, denoted as $\delta_{c}$.
-Specifically, $\delta_{c}$ is inversely related to the highest validation score achieved over all model parameters $\theta$:
+To address the challenges posed by high task diversity, we propose a novel approach inspired by the curriculum learning paradigm to improve both the convergence speed and the generalization ability of the meta-learner.
+This approach adopts a sequential task presentation strategy, following the principle of “easier tasks first, harder tasks later.” To implement this, we introduce a teacher recommender, which undergoes initial training for each auxiliary task and utilizes the highest validation score to estimate task difficulty.
+This difficulty metric guides the curriculum design for task sampling during each meta-training iteration.
+Specifically, to measure task difficulty, we split the samples into a training set $D_{\text{train}}$ and a validation set $D_{\text{valid}}$.
+A base LSTM model is trained extensively on $D_{\text{train}}$ to serve as the teacher recommender.
+The difficulty of a task is quantified using a performance metric $\Phi$ (e.g., AUC) evaluated on $D_{\text{valid}}$.
+The task difficulty, denoted as $\delta_{c}$, is inversely proportional to the highest validation score obtained across all model parameters $\theta$:
 $\delta_{c}\propto-\max_{\theta}\Phi(f_{\theta},D_{\text{valid}_{c}}|D_{\text{train}_{c}})$ (6)
 Intuitively, a lower peak validation score indicates more challenging vibration signal pattern within a task.
-II-B Target-level Fine-tune
-After obtaining the general pattern of auxiliary tasks through the meta-learning paradigm, it is important to further fine-tune the learning model for the target task.
-By investigating the freezing of layers and then fine-tuning, the network generalizes better than one trained directly on the target dataset [24].
-This operation not only incorporates information from the auxiliary tasks but also allows for better adaptation to the target task [21].
-Specifically, assuming the model (e.g., LSTM) through meta-level training contains $L$ layers, we freeze first $l$ $(1\leq l\leq L)$ layers, while adding $n$ layers after the $l$ layers and fine tune the parameters by using data of the target task, so as to maximize the reusability of the general parameters.
-Hence, $\mathbf{e}_{hist}^{(tar)}$ is fed into the recurrent layer to infer the hidden state $\boldsymbol{h^{u}_{t_{k}}}$ at $t_{k}$, given by,
-$\boldsymbol{h^{u}_{t_{k}}}=LSTM_{frozen}(\mathbf{e}_{hist}^{(tar)}).$ (7)
-By doing this, the freezing-layer operation helps generate a network, capable of better balancing the parameters between the auxiliary tasks and target task after the final model fine-tuning.
-The probability distribution on all working conditions is calculated by softmax function:
-$\boldsymbol{\hat{y}}=softmax(f(\mathbf{h}^{u}_{t_{k}}))$ (8)
-where $f$ is a fully connected layer to transform $\boldsymbol{h^{u}_{t_{k}}}$ into a $|\mathcal{P}|$-dimensional vector, and $|\mathcal{P}|$ is the number of faulty class in the target task.
-Hence, the objective function is defined by:
-$\mathcal{J}=-\sum_{i=1}^{|\mathcal{P}|}\mathbf{y}[i]\cdot log(\hat{\mathbf{y}}[i])$ (9)
-where $\mathbf{y}$ is an one-hot embedding of the ground-truth fault.
-Algorithm 1 summarizes the training process of CMEFD, consisting of meta training (lines 3-11), freezing layers and model fine tuning (lines 12-14), as well as fault diagnosis (lines 15-16).
 Algorithm 1 Related Task Aware Curriculum Meta-learning (RT-ACM) Enhanced Fault Diagnosis framework
@@
 8:        Calculate the gradient update of $\theta^{\prime}_{y_{m}}$ by Eq.(5);
-9:        Sample another $N$ vibration signal from $\mathbb{D}_{i}$ as the eval_batch based on the difficulty;
+9:        Sample another $N$ signal from $\mathbb{D}_{i}$ as the eval_batch based on the difficulty;
 10:     end for
@@
 16:  Calculate the prediction loss for each record via Eq.(9);
+II-B Target-level Fine-tune
+After obtaining the general pattern of auxiliary tasks through the meta-learning paradigm, it is important to further fine-tune the learning model for the target task.
+By investigating the freezing of layers and then fine-tuning, the network generalizes better than one trained directly on the target dataset [28].
+This operation not only incorporates information from the auxiliary tasks but also allows for better adaptation to the target task [29].
+After extracting general patterns from auxiliary tasks via the meta-learning paradigm, fine-tuning the model for the target task becomes essential [28].
+By investigating the freezing of layers and then fine-tuning, the network generalizes better than one trained directly on the target dataset.
+This approach not only integrates knowledge from auxiliary tasks but also facilitates better adaptation to the target task [29].
+Assuming the meta-trained model (e.g., LSTM) consists of $L$ layers, we freeze the first $l$ layers $(1\leq l\leq L)$, add $n$ new layers after the frozen ones, and fine-tune the parameters using target task data to maximize the reuse of general parameters.
+Consequently, the target task input $\mathbf{e}_{hist}^{(tar)}$ is passed through the recurrent layer to compute the hidden state $\boldsymbol{h^{u}_{t_{k}}}$ at time $t_{k}$, as expressed by:
+$\boldsymbol{h^{u}_{t_{k}}}=LSTM_{frozen}(\mathbf{e}_{hist}^{(tar)}).$ (7)
+By doing this, the freezing-layer operation facilitates the creation of a network that achieves a more effective balance between the parameters learned from auxiliary tasks and those optimized for the target task, following the final fine-tuning of the model.
+The probability distribution on all working conditions is calculated by softmax function:
+$\boldsymbol{\hat{y}}=softmax(f(\mathbf{h}^{u}_{t_{k}}))$ (8)
+where $f$ is a fully connected layer to transform $\boldsymbol{h^{u}_{t_{k}}}$ into a $|\mathcal{P}|$-dimensional vector, and $|\mathcal{P}|$ is the number of faulty class in the target task.
+Hence, the objective function is defined by:
+$\mathcal{J}=-\sum_{i=1}^{|\mathcal{P}|}\mathbf{y}[i]\cdot log(\hat{\mathbf{y}}[i])$ (9)
+where $\mathbf{y}$ is an one-hot embedding of the ground-truth fault.
+Algorithm 1 summarizes the training process of RT-ACM, consisting of meta training (lines 4-12), freezing layers and model fine tuning (lines 13-14), as well as fault diagnosis (lines 15-16).
 III Case Study
-TABLE I: COMPARISON OF TRADITIONAL DL-BASED AND META-LEARNING-BASED MODELS
-(*The best results are highlighted in bold.)
+TABLE I: COMPARISON OF DL-BASED AND META-LEARNING-BASED MODELS
+(*The best results are highlighted in bold; the runner up is underlined.)
 3-way 6-way
@@
 T1 T2 T3 T4 T1 T2 T3 T4 T1 T2 T3 T4 T1 T2 T3 T4
-Deep Learning Based
-LSTM 34.65 32.80 37.80 40.75 52.73 42.97 46.46 51.95 20.57 18.32 25.71 27.95 32.86 27.45 30.99 34.67
 Cap-Net 41.23 35.56 45.24 47.89 69.53 55.43 65.31 63.74 37.53 31.83 41.02 44.65 63.18 50.29 59.55 57.38
 WDCNN 41.50 34.48 45.90 46.07 69.84 55.75 66.93 63.23 37.82 30.92 42.45 44.03 63.68 50.53 60.29 57.05
-Meta Learning Based
 MAML 85.23 80.92 83.95 85.80 97.71 95.72 97.03 96.75 70.25 66.38 69.22 71.45 73.59 70.87 72.13 71.68
-WCFS 96.77 95.52 97.03 97.70 98.70 97.59 98.88 98.73 85.67 83.32 89.15 87.49 89.68 87.97 93.27 90.15
-CHAML 92.95 90.64 93.22 94.86 95.00 93.50 96.21 96.00 88.41 86.61 88.31 90.17 92.67 91.29 90.16 93.61
+TMSL 97.21 96.90 97.87 97.80 98.20 97.50 99.47 99.66 86.00 82.18 88.71 89.32 89.96 87.29 92.13 91.45
 RT-ACM 98.51 97.44 98.62 98.53 99.68 99.53 99.73 99.56 93.22 92.49 94.38 95.14 97.25 96.59 97.48 97.05
-Rolling bearings play a vital role as essential components in electric machines and find extensive usage across various mechanical equipment.
-Given the prolonged rotational wear and exposure to high temperatures during operation, bearing failures are common faults that occur.
-Therefore, conducting research on fault diagnosis in rolling bearings holds significant importance.
-This study specifically focuses on the detection of faults in rolling bearings within electric motor, considering the varying working conditions and a limited number of available samples.
-To evaluate the effectiveness of the proposed model, a series of case studies are conducted using two widely recognized datasets: the Case Western Reserve University (CWRU) Bearing Dataset [25], the Paderborn University Rolling Bearing Dataset [26] and the Gearbox dataset [19].
-Moreover, several state-of-the-art models have been applied for comparison, including LSTM [23], the capsule network (Cap-Net) [4], wide deep CNN (WDCNN) [27], MAML [17], WDCNN-few-shot (WCFS) [10], CHAML [20], PKAML [19].
-The LSTM, Cap-Net and WDCNN are DL-based model, MAML, WCFS, CHAML and PKAML are meta-learning based models.
-(a) Healthy
-(b) Ball 0.007
-(c) Inner Race 0.007
-(d) Outer Race 0.007
-Figure 2: The visualisation of the vibration signals corresponding to different fault types under same fault diameter and working condition.
-III-A Case 1: CWRU Bearing Dataset
-III-A1 Dataset Description
-The CWRU dataset has been commonly used to verify DL fault diagnosis methods.
-The selected rolling CWRU bearing dataset encompasses a comprehensive range of fault types, as depicted in Table II.
-The dataset comprises four distinct types of electric machine health states, which encompass one normal state and three fault states.
-The sizes of the faults are specified in terms of diameters, measuring 0.007, 0.014, and 0.021 inches.
-Fig.
-2 provides a visual representation of the vibration signals corresponding to different fault types, all with the same fault diameter.
-As mentioned in Section II, fault diagnosis under various working conditions can be regarded as different tasks, and each task is a multi-class classification problem.
-Detailed descriptions of each task and relationship between the motor load and motor speed shown are presented in Table III.
-We treat one of the tasks as the target task and the rest as auxiliary tasks each time.
-Following [28], we chronologically divide the dataset of the target task into training, validation, and test sets with a ratio of 8:1:1.
+Rolling bearings are critical components in electric machines and are widely utilized in various mechanical systems.
+Due to prolonged rotational stress and exposure to high temperatures, bearing failures frequently occur during operation.
+As a result, research on fault diagnosis for rolling bearings is of paramount importance.
+This study concentrates on detecting faults in rolling bearings within electric motors, taking into account diverse working conditions and the limited availability of samples.
+To assess the performance of the proposed model, a series of case studies are carried out using two well-known datasets: the Case Western Reserve University (CWRU) Bearing Dataset [30] and the Paderborn University Rolling Bearing Dataset [31].
+Additionally, several state-of-the-art models are employed for comparison, including the capsule network (Cap-Net) [6], wide deep CNN (WDCNN) [32], MAML, TMSL, CHAML, and PKAML.
+Among these, Cap-Net and WDCNN are deep learning-based models, while MAML, CHAML, and PKAML belong to the category of meta-learning-based approaches.
 TABLE II: INFORMATION OF USED BEARINGS FROM CWRU DATASET
@@
 $T^{4}$ 3 10 1730
-III-A2 Hyper-parameter Settings
-The optimal hyper-parameter settings for all models are empirically found out based on the performance on the validation set.
-Specifically, for TL-based moded, the learning rate is selected from {0.1, 0.05, 0.01, 0.005, 0.001, 0.0001}, and batch size is set from {256, 128, 64}.
-For meta-learning based models, the learning rate $\alpha,\beta$ are searched from {0.1, 0.01, 0.001, 0.0001}; and the batch size is set from {256, 128, 64} for a fair comparison.
-For RT-ACM, the number of freezing laysers $l$ is searched in range of [1,4] stepped by one, where the best setting is 3 for all tasks.
-Figure 3: Accuracy in the (a) outer loop and (b) inner loop.
-Figure 4: Model performance on $T^{1}$.
-III-A3 Performance Comparison
-Table I presents a comprehensive comparison of traditional deep learning-based models and meta-learning-based models across various tasks and shots.
+III-A Case A: CWRU Bearing Dataset
+III-A1 Dataset Description
+The CWRU dataset is widely utilized to evaluate fault diagnosis methods [2, 15, 33].
+The selected rolling bearing dataset from CWRU includes a comprehensive set of fault types, as shown in Table II.
+It covers four distinct electric machine health states, including one normal state and three fault states.
+The fault sizes are specified by their diameters, which measure 0.007, 0.014, and 0.021 inches.
+As outlined in Section II, fault diagnosis under different working conditions can be considered as separate tasks, with each task representing a multi-class classification problem.
+Table III provides detailed descriptions of each task, along with the relationship between motor load and motor speed.
+For each iteration, one of the tasks is treated as the target task, while the remaining tasks are used as auxiliary tasks.
+Following [34], we chronologically divide the dataset of the target task into training, validation, and test sets with a ratio of 8:1:1.
+III-A2 Performance Comparison
+Table I presents a comprehensive comparison of deep learning-based models and meta-learning-based models across various tasks and shots.
 The results indicate a clear performance superiority of meta-learning-based models over their traditional counterparts.
 Specifically, our method RT-ACM consistently achieves the highest accuracy across all scenarios, notably outperforming other models with an impressive accuracy range of 98.51% to 99.73% in the 3-way tasks and 93.22% to 97.48% in the 6-way tasks.
-MAML, WCFS and CHAML, while also meta-learning-based, demonstrate strong performances, especially in the 3-way 5-shot settings.
-CHAML generally outperforms WCFS due to its structured learning process.
-By leveraging a curriculum sampling, CHAML progressively exposes the model to increasingly complex tasks, enhancing its ability to generalize and perform well across diverse scenarios.
-This staged learning process allows the model to build a more robust understanding of the task at hand, resulting in higher accuracy, particularly in more challenging 6-way tasks.
+MAML and TMSL, while also meta-learning-based, demonstrate strong performances, especially in the 3-way 5-shot settings.
+Fig.
+3(a) presents the 3-way 5-shot performance results on Task 1, while Fig.
+3(b) illustrates the 6-way 5-shot performance on the same target task.
+It is evident that TMSL generally outperforms MAML, attributed to its advantage in sequence learning processes.
 Our proposed method RT-ACM integrates related task-aware meta-learning with a curriculum strategy, which further amplifies its performance.
@@
 Fig.
-3a shows the performance comparison of four meta-learning-based models, in terms of accuracy over increasing global update steps.
+4(a) shows the performance comparison of four meta-learning-based models, in terms of accuracy over increasing global update steps.
 Our proposed method, RT-ACM, which leverages related task-aware meta-learning and a curriculum strategy, demonstrates a superior ability to rapidly achieve high accuracy.
 This rapid improvement can be attributed to the efficient transfer of knowledge from related tasks and a structured learning process that progresses from simpler to more complex tasks.
-In contrast, CHAML, which employs curriculum sampling, lags behind RT-ACM and fails to match its performance throughout the training process due to a less effective related task prioritization.
-WCFS initially shows competitive accuracy but experiences a decline in performance in later stages, possibly due to overfitting or poor generalization.
+In contrast, TMSL, which employs k-nn based task evaluation, lags behind RT-ACM and fails to match its performance throughout the training process due to a less effective related task prioritization.
 MAML, while maintaining steady progress, does not achieve the same level of accuracy as RT-ACM.
+WDCNN initially shows competitive accuracy but experiences a decline in performance in next stages, possibly due to overfitting or poor generalization.
 Fig.
-3b illustrates the relationship between the performance of model and the number of local update steps.
-Through empirical analysis, it is observed that a single-step update is adequate to enhance the accuracy significantly.
-Fig.
-4 illustrates that the RT-ACM model stands out for its exceptional combination of efficiency and accuracy on 3-way 5-shot.
-When compared to conventional models like LSTM and WDCNN, RT-ACM’s testing time is considerably shorter,which reflecting its higher processing speed.
-Additionally, RT-ACM competes closely in accuracy with more sophisticated models like MAML and WCFS.
-These results highlight the potential of RT-ACM, underscoring the advantages of incorporating auxiliary tasks to enhance the target task in terms of accuracy.
-Moreover, the curriculum transfer of knowledge from auxiliary tasks in RT-ACM demonstrates rapid learning capabilities.
-(a) LSTM
-(b) MAML
-(c) WCFS
-(d) RT-ACM
-Figure 5: The visualization of classification from four models.
-(a) LSTM
-(b) MAML
-(c) WCFS
-(d) RT-ACM
-Figure 6: The visualization of confusion matrices from four models.
-III-B Case 2: PU Bearing Dataset
+4(b) demonstrates the superior performance of the RT-ACM model in both running time in each iteration and accuracy on the 3-way 5-shot task.
+While RT-ACM requires slightly more computational time compared to deep learning models like Cap-net and WDCNN due to its meta-learning process of acquiring learning capabilities through different tasks, it significantly reduces the time consumption compared to other meta-learning approaches like MAML and TMSL.
+This improvement in efficiency can be attributed to the curriculum strategy’s progressive learning approach, moving from simple to complex tasks.
+Furthermore, RT-ACM achieves competitive accuracy levels, highlighting its potential in balancing the trade-off between computational efficiency and model performance.
+The incorporation of auxiliary tasks has proven beneficial in improving accuracy on the target task.
+Figure 3: Comparison of (a) 3-Way and (b) 6-Way 5-Shot Performance on $T_{1}$.
+Figure 4: Accuracy in the (a) outer loop and (b) model performance on $T^{1}$.
+III-B Case B: PU Bearing Dataset
 III-B1 Dataset Description
@@
 To mitigate the influence of similar bearings on the dataset, vibration signals were collected from 26 bearings characterized by artificial damage, natural damage, and healthy states.
-The detailed working conditions are presented in Table IV.
-Following [19], the first 20 bearings were utilized for training, while the remaining 6 bearings were allocated for testing.
-Each bearing exhibited four distinct working conditions are presented in Table V, primarily varying in rotational speed, load torque, and radial force.
-TABLE IV: INFORMATION OF USED BEARINGS FROM PU DATASET
-No. Bearing Code Faulty Type Damage Method
+The detailed working conditions are presented in Table V.
+Following [8], 32 unique bearing fault types are evaluated across four distinct operating conditions.
+To reflect the complexity of real-world industrial environments and to enhance the precision of diagnostic outcomes, each fault type is considered a separate category under each working condition, resulting in a total of 128 fine-grained fault categories.
+From these, 100 categories are randomly chosen for training purposes, while the remaining 28 are reserved for testing.
+Further specifics of the working conditions are provided in Table VI.
+TABLE IV: THE COMPARATIVE METHODS IN PU DATASET ($T^{1}$)
+(*The best results are highlighted in bold; the runner up is underlined.)
+Models Accuracy Precision F1-score
+5-way
+5-shot 6-shot 5-shot 6shot 5-shot 6-shot
+Cap-Net 46.61 48.85 46.52 48.03 46.36 48.47
+WDCNN 72.48 74.06 71.46 74.57 68.47 75.08
+MAML 89.65 90.9 87.90 88.92 89.06 90.77
+[…]
+RTACM 94.34 95.65 94.21 95.42 94.27 96.18
+TABLE V: Information of Bearings Used from PU Dataset
+No. Bearing Code Fault Type Damage Method
 1 K003 Healthy -
-2 K004 Helathy -
+2 K004 Healthy -
 3 K005 Healthy -
 4 K006 Healthy -
-5 KA01 Outer Race Muliple Damage
+5 KA01 Outer Race Multiple Damage
 6 KA03 Outer Race Electric Engraver
@@
 9 KA08 Outer Race Drilling
-10 KI01 Inner Race Muliple Damage
+10 KA09 Outer Race Drilling
+11 KI01 Inner Race Multiple Damage
+12 KI03 Inner Race Electric Engraver
 13 KI05 Inner Race Electric Engraver
 14 KI07 Inner Race Electric Engraver
+15 KI08 Inner Race Electric Engraver
+16 KA04 Outer Race Fatigue
 17 KA15 Outer Race Plastic Deform
@@
 19 KA22 Outer Race Fatigue
+20 KA30 Outer Race Plastic Deform
 21 KB23 Inner/Outer Fatigue
 22 KB27 Inner/Outer Plastic Deform
+23 KI04 Inner Race Fatigue
 24 KI16 Inner Race Fatigue
@@
 32 KI18 Inner Race Fatigue
-TABLE V: DETAILED DESCRIPTIONS OF EACH TASK UNDER PU DATASET
+TABLE VI: DETAILED DESCRIPTIONS OF EACH TASK UNDER PU DATASET
 Task Rational speed (rpm) Load torque (Nm) Radial force (N)
@@
 $T^{4}$ 1500 0.7 1000
-TABLE VI: THE COMPARATIVE METHODS IN PU DATASET ($T^{1}$)
-(*The best results are highlighted in bold.)
-Models Accuracy Precision F1-score
-10-way
-1-shot 5-shot 1-shot 5-shot 1-shot 5-shot
-LSTM 40.15 50.89 39.92 50.05 39.61 49.84
-MAML 76.85 85.35 76.71 85.19 76.63 85.12
-WCFS 78.52 91.07 78.37 90.92 78.45 91.03
-CHAML 78.75 90.20 77.61 91.07 78.58 91.13
-PKAML 79.20 92.09 79.13 91.97 79.16 92.01
-RT-ACM 81.57 94.34 81.43 94.21 81.46 94.27
-(a) Training Loss
-(b) Validation Accuracy
-Figure 7: The visualization of training loss and validation accuracy from four models.
-III-B2 Evaluation Metrics
-Given the inclusion of artificially damaged data in the PU dataset, resulting in imbalanced distribution, precision and F1 score offer better insights into model performance.
-Thus, three established evaluation metrics are employed [6]: Accuracy, F1-Score and Precision.
-Accuracy quantifies the proportion of correct predictions in the total predictions made.
-It is given by the formula:
-Accuracy=Number of Correct PredictionsTotal Number of $\text{Accuracy}=\frac{\text{Number of Correct Predictions}}{\text{Total Numberof Predictions}}$
-Accuracy provides a general measure of model performance, especially when the class distributions are balanced.
-Precision is the fraction of relevant instances among the retrieved instances.
-In other words, it quantifies the accuracy of positive predictions.
-It is given by:
-Precision=True PositivesTrue Positives+False $\text{Precision}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Positives}}$
-Precision provides insight into the reliability of positive predictions made by the model.
-The F1 score is the harmonic mean of precision and recall.
-It balances the trade-off between precision and recall and is especially useful when the class distribution is imbalanced.
-It is computed as:
-[…]
-Where, precision is the proportion of true positive predictions in the set of positive predictions made. Recall is the proportion of true positive predictions in the set of actual positives.
-III-B3 Performance Comparison
-Table VI shows a performance comparison of various models on the PU dataset in 10-way, 1-shot and 5-shot.
-Notably, all methods except LSTM are meta-learning based approaches, allowing them to handle data sparsity issues more effectively than LSTM..
-Among these methods, PKAML employs a pretraining process to integrate prior knowledge, enhancing its performance.
-CHAML utilizes a curriculum sampling in meta-training, providing a structured approach to learning.
-However, both PKAML and CHAML do not adequately consider the diversity and relevance of auxiliary tasks in relation to the target task, which can limit their effectiveness.
+III-B2 Performance Comparison
+Table IV shows a performance comparison of various models on the PU dataset in 5-way, 5-shot and 6-shot.
+Notably, all methods except Cap-Net and WDCNN are meta-learning based approaches, allowing them to handle data sparsity issues more effectively than traditional DL-based model.
+Among these methods, MAML employs a meta-training process to integrate different tasks, enhancing its performance.
+TMSL utilizes a k-nn based sequence sampling in meta-training, providing a structured approach to learning.
+However, both MAML and TMSL do not adequately consider the diversity and relevance of auxiliary tasks in relation to the target task, which can limit their effectiveness.
 In contrast, our proposed method, RT-ACM, incorporates related task-aware meta-learning along with a curriculum strategy, resulting in superior performance.
-This is evidenced by RT-ACM achieving the highest scores across all metrics and shots, with an accuracy of 81.57% and 94.34% in 1-shot and 5-shot settings, respectively, outperforming the runner-up PKAML and other models significantly.
+This is evidenced by RT-ACM achieving the highest scores across all metrics and shots, with an accuracy of 94.34% and 95.65% in 5-shot and 6-shot settings, respectively, outperforming the runner-up TMSL and other models significantly.
 Fig.
 5 presents t-SNE visualizations for four models, highlighting their ability to discriminate between classes.
-The LSTM visualization shows poorly separated clusters, indicating significant class overlap and lower classification efficacy.
+The Cap-Net visualization shows poorly separated clusters, indicating significant class overlap and lower classification efficacy.
 MAML improves with tighter clusters but still suffers from overlap at the cluster centers.
-WCFS slightly betters class separation, with clearer clusters but some outliers.
+TMSL slightly betters class separation, with clearer clusters but some outliers.
 RT-ACM’s visualization stands out, with each class forming distinct, isolated clusters, showcasing superior class discrimination and robust classification.
@@
 6 showcases confusion matrices for four classification model, illustrating each model’s predictive performance.
-The LSTM matrix shows widespread misclassifications, such as confusing class 2 with class 3, indicating difficulty in distinguishing similar classes.
-MAML improves, concentrating predictions along the diagonal but still mixes up classes like 4 and 5.
-WCFS further enhances accuracy, with clearer diagonal values and fewer errors, notably between classes 7 and 8.
-RT-ACM stands out with a strong diagonal focus, demonstrating superior accuracy and minimal misclassification, especially distinguishing between closely related classes like 5 and 6, and 9 and 10, where other models struggle.
+The Cap-Net matrix shows widespread misclassifications, such as confusing class 3 with class 4, indicating difficulty in distinguishing similar classes.
+MAML improves, concentrating predictions along the diagonal but still mixes up classes like 4 and 9.
+TMSL further enhances accuracy, with clearer diagonal values and fewer errors, notably between classes 9 and 10.
+RT-ACM stands out with a strong diagonal focus, demonstrating superior accuracy and minimal misclassification, especially distinguishing between closely related classes like 4 and 10, where other models struggle.
 This highlights RT-ACM’s robust predictive capability across diverse classes.
+Figure 5: The visualization of classification from four models on (10-way 5-shot).
+Figure 6: The visualization of confusion matrices from four models (10-way 5-shot).
+Figure 7: (a) The performance of models across varying numbers of local update steps and (b) the impact of the number of freezing layers.
 Fig.
-7a illustrates the training loss of various models, including LSTM, MAML, WCFS, and RT-ACM.
-The LSTM exhibits significant fluctuations, indicating instability from a small dataset size and potential overfitting issues.
-In contrast, RT-ACM shows a quick and steady reduction in loss, outperforming MAML and WCFS in terms of learning efficiency and model regularization, suggesting a superior dataset fit.
-Fig.
-7b compares validation accuracy, with RT-ACM achieving higher accuracy early and maintaining it, highlighting its robustness and generalization capability.
-LSTM’s accuracy fluctuates, while MAML and WCFS improve steadily but don’t match RT-ACM’s performance.
-RT-ACM excels in convergence speed and generalization, benefiting from task-aware meta-learning and a curriculum strategy.
+7(a) illustrates a sensitivity analysis of a meta-learning-based model across varying numbers of local update steps.
+The results indicate that a single-step update is sufficient to significantly enhance accuracy, demonstrating that MAML-based models can achieve notable improvements with minimal updates.
+In Fig.
+7(b), the impact of the number of frozen layers on model performance is examined.
+The performance initially improves as the number of frozen layers increases, reaching a peak when three layers are frozen, after which performance slightly declines.
+This suggests that three frozen layers offer the optimal balance for performance.
+Too few frozen layers may lead to a loss of auxiliary information, while too many frozen layers limit the model’s ability to adapt to the target task, resulting in overfitting.
 IV Conclusion and Future Work
-In this paper, we develop the Related Task Aware Curriculum Meta-learning (RT-ACM) enhanced fault diagnosis framework, a novel approach in the realm of fault diagnosis for smart manufacturing designed to address the issue of data scarcity.
-RT-ACM innovatively utilizes auxiliary tasks to enrich the primary fault diagnosis tasks, adhering to the principle of “paying more attention to more relevant knowledge” by focusing on the most relevant knowledge and adopting an “easy-to-hard” curriculum learning strategy.
-Our comprehensive evaluations across three real-world datasets confirm the effectiveness and superiority of the RT-ACM framework.
+In this paper, we introduced a Related Task Aware Curriculum Meta-learning (RT-ACM) enhanced fault diagnosis framework, a novel approach in the realm of fault diagnosis for smart manufacturing designed to address the issue of data scarcity.
+RT-ACM introduces an innovative approach by employing related task-aware meta-learning to enhance primary fault diagnosis tasks.
+It adheres to the principle of “paying more attention on relevant knowledge” by prioritizing the most pertinent information and following an “easy-to-hard” curriculum strategy.
+Our comprehensive evaluations across two real-world datasets confirm the effectiveness and superiority of the RT-ACM framework.
 The experiments not only highlight the efficacy of the framework but also validate the architectural choices, underscoring the potential of combining curriculum strategy with related task aware meta-learning for enhanced fault diagnosis.